OpenStudy (laddiusmaximus):

lim---->0 4.5^(2x)-2.25/4.5^x-1.5

OpenStudy (anonymous):

replace x by zero, see what you get

OpenStudy (laddiusmaximus):

I did that but the answer wasnt right.

OpenStudy (anonymous):

i get \[\frac{1-2.25}{1-1.5}=\frac{1.25}{.5}=2.5\] does that look right?

OpenStudy (laddiusmaximus):

yeah. I figured you had to use a number a little more than zero. I had tried 0.1 but I didnt go far enough. thank you